# Penalized Fair Regression for Multiple Groups in Chronic Kidney Disease

### v1

@@ Abstract
Fair regression methods have the potential to mitigate societal bias concerns in health care, but there has been little work on penalized fair regression when multiple groups experience such bias. We propose a general regression framework that addresses this gap with unfairness penalties for multiple groups. Our approach is demonstrated for binary outcomes with true positive rate disparity penalties. It can be efficiently implemented through reduction to a cost-sensitive classification problem. We additionally introduce novel score functions for automatically selecting penalty weights. Our penalized fair regression methods are empirically studied in simulations, where they achieve a fairness-accuracy frontier beyond that of existing comparison methods. Finally, we apply these methods to a national multi-site primary care study of chronic kidney disease to develop a fair classifier for end-stage renal disease. There we find substantial improvements in fairness for multiple race and ethnicity groups who experience societal bias in the health care system without any appreciable loss in overall fit.
Keywords: fair regression, penalized regression, algorithmic bias, chronic kidney disease
@@ 1 INTRODUCTION
Fair regression methods modify the loss function to encode fairness measures. These loss functions typically include a single penalty or constraint based on a measure of unfairness for one group (calders_controlling_2013; berk_convex_2017; corbett-davies_measure_2023). Subsequent work has developed additional technical methods for solving single group penalized or constrained fair regression problems (dwork_decoupled_2017; chzhen_fair_2020; lohaus_too_2020; mishler_fade_2021) and tailored them for applications in cyberbullying (gencoglu_cyberbullying_2021) and health policy (zink_fair_2020).
@@
Although there is a mathematical equivalence between penalized and constrained methods such that each penalized regression has a corresponding constrained regression with the same solution, most methods formulate their fair regressions as constrained optimization problems. This overlooks that penalty-based approaches have several advantages relative to constraint-based approaches that warrant further development. For one, the approaches are conceptually different. A constraint is a user-imposed maximum level of allowable unfairness, while a penalty is a user-imposed tradeoff between accuracy and fairness. These conceptions are appropriate for different applications and their different social and regulatory contexts. The two approaches may also require different computational implementations. Much of the constrained optimization literature involves development of original solution algorithms (kearns_preventing_2018; komiyama_nonconvex_2018), whereas there are many existing algorithms for minimizing penalized loss functions under different conditions, including gradient descent, Newton’s method, Nelder-Mead, and their variations (nelder_simplex_1965; shanno_conditioning_1970; loshchilov_decoupled_2019). Outside of the fairness literature, penalized methods are commonly used in statistics applications (hoerl_ridge_1970; tibshirani_regression_1996) and fair penalty methods can benefit from advances established in those applications.
@@
Fair regression has the potential to reduce the impact of societal bias in health care. However, only a small number of fair regression works have examined realistic use cases in health care (zink_fair_2020; mcguire_improving_2021; pfohl_empirical_2021; pfohl_net_2022; wang_enhancing_2025). Health care applications have complex ethical considerations, many of which are unique due to the complexity of medical systems (chen_ethical_2021; chen_algorithmic_2023; feng_fair_2025). For example, questions about how to handle certain patient characteristics in clinical algorithms, such as race, are nuanced such that not all fair regression approaches may be appropriate (cusick_algorithmic_2024; zink_race_2024).
@@
A major limitation of most fair regression approaches is that they focus on fairness with respect to a single group when multiple groups typically face unfairness in real world settings (yang_fairness_2020; chen_ethical_2021; liu_practical_2022; wang_enhancing_2025). The joint consideration of multiple groups in fair regression is non-trivial, with challenges related to the optimal weighting of different groups as well as the additional computational complexity (kearns_preventing_2018; narasimhan_approximate_2020; diana_minimax_2021). We refer here to multiple groups rather than sensitive attributes to avoid the particular social and sometimes legal connotations of the latter, reflecting the breadth of possible groups for which these methods may be appropriate (deho_should_2023).
@@
While limited compared to methods for single groups, there is a small literature on fair constrained regression for multiple groups establishing basic frameworks and efficient solution methods (kearns_preventing_2018; narasimhan_learning_2018; agarwal_fair_2019; zafar_fairness_2019; narasimhan_approximate_2020; yang_fairness_2020; diana_minimax_2021; hu_sequentially_2024; wang_enhancing_2025). Of particular note, agarwal_reductions_2018 proposed a method that reduces fair binary classification to a sequence of cost-sensitive classification problems, enabling the use of efficient off-the-shelf classifiers for fair constrained regression. This approach is flexible across common fairness metrics and was extended in multiple later papers, including to continuous outcomes (agarwal_fair_2019) and to non-linear but convex constraints, including demographic parity and equalized odds, and convex or fractional-convex losses (narasimhan_learning_2018). Much of the constrained optimization literature involves solving a min–max saddle-point game between a predictor and an adversary enforcing fairness constraints, often leading to non-convex objectives, instability, substantial hyperparameter sensitivity, and high computational cost, especially when group sizes are imbalanced.
@@
In contrast, we have identified even fewer papers that implement fair penalized regression with respect to multiple groups, further highlighting the lesser focus on this formulation of the estimation problem. perez-suay_fair_2017 used the Hilbert Schmidt independence criterion for fairness penalties. This method is only compatible with one measure of fairness, demographic parity, which is satisfied when estimated outcomes are independent of groups. A ridge penalty method that builds on komiyama_nonconvex_2018 to enforce demographic parity was developed in scutari_achieving_2022. Their approach notably includes group identifiers as predictors, which will not be appropriate for some estimation problems. They also proposed an extension to other fairness metrics, including equalized odds. The focus of these first two papers on demographic parity restricts applications in health care, where the most pressing fairness concern may differ from and be incompatible with demographic parity (seyyed-kalantari_underdiagnosis_2021; chen_algorithmic_2023). pfohl_empirical_2021 and pfohl_net_2022 implement fair penalized regression with a single penalty term that measures unfairness with respect to a single categorical group, capturing a subset of multiple group fair regression problems. Lastly, liu_practical_2022 introduced a fairness measure built on cross-covariance operators on reproducing kernel Hilbert Spaces which were used to construct unfairness penalties. While the approach is flexible with respect to fairness measures, its complexity presents a potential barrier to its implementation and no code is available.
@@
Our application creates a fair penalized regressor classifying whether high-risk individuals with chronic kidney disease (CKD) will develop end stage renal disease (ESRD) in the next year. CKD is prevalent, affecting an estimated 14% of the US adult population (cdc_chronic_2025). In severe cases, CKD progresses to ESRD, requiring either dialysis or kidney transplant, but progress can be delayed through treatment and management. Black, Hispanic, Native American, and Asian patients all have higher incidence and prevalence of ESRD than white patients, and they have worse kidney function, on average, than white patients when initiating treatment for ESRD (johansen_us_2025). Identifying patients at risk for ESRD could help these patients engage in more intensive treatment and slow CKD progression, and fair regression is a potential approach to partially address disparities in progression to ESRD. Another key algorithm for defining CKD progression, the estimated glomerular filtration rate (eGFR) equation, was recently revised to remove race adjustment, so CKD may be a context where it is especially appropriate to use a fair regression method that can increase fairness without using race adjustments (cusick_algorithmic_2024).
@@ 2.1 Fairness Formulation
We use true positive rate equality, based on equal opportunity (hardt_equality_2016), which holds if the true positive rate of our estimator for $Y$ is the same for members of group $g$ as it is for a reference group $r$. A focus on true positive rate fairness (or, equivalently, false negative rate fairness) is appropriate for the problem of low rates of screening and diagnosis among high-risk patients in certain groups (seyyed-kalantari_underdiagnosis_2021). This is especially true in cases where the benefit of a diagnosis is much greater than the potential harm, as this implies that false negatives are much more costly than false positives and should be the primary fairness concern. We can express different versions of this fairness measure for estimated outcomes and for estimated probabilities. The estimated outcome formulation is more interpretable. However, the estimated probability formulation is differentiable, so it can be easily minimized, and it does not depend on a particular classification threshold.
@@
where $\hat{Y}$ represents predicted outcomes. For estimated probabilities, we substitute $\hat{P}$ for $\hat{Y}$ as has been done in prior literature (wang_enhancing_2025). This version of the unfairness measure, which we call $U_{EOP}$, is given by:
@@ 2.2 Fairness Measures for Multiple Groups
There is no consensus in the literature on optimal ways to evaluate fairness for multiple groups simultaneously, as much of the past fair regression work characterizes unfairness for each group individually (agarwal_reductions_2018; komiyama_nonconvex_2018). Given the unfairness metrics in Equation 1 and Equation 2 for a given group and reference, which we will denote $U_{g}$, we study three different approaches to aggregating across groups.
@@
    Additionally, we study a maximum across all group-level unfairness measures, based on ghosh_characterizing_2022: $U_{Max}=\max_{g\in G}U_{g}$.
@@ 3 METHODS
Our novel penalized fair regression approach generalizes to different fairness metrics, outcome variable types, and estimator architectures. We focus our main demonstration of the method on logistic regression with true positive rate disparity penalties. The online appendix contains loss functions for continuous outcomes with the mean residual difference as the unfairness metric.
@@ 3.1 Regression with Fairness Penalties
$L=-\sum_{i=1}^{n}(Y_{i}\ln(\hat{P}_{i})+(1-Y_{i})\ln(1-\hat{P}_{i}))+\sum_{g\in G}\lambda_{g}n_{g}\left(\frac{\sum_{k\in r}Y_{k}\hat{P}_{k}}{\sum_{k\in r}Y_{k}}-\frac{\sum_{j\in g}Y_{j}\hat{P}_{j}}{\sum_{j\in g}Y_{j}}\right).$
The above loss function is convex but does not have a closed form minimizing solution. We can calculate the gradient and use a gradient descent-based algorithm to find the solution. With $\nabla\hat{P}=\hat{P}(1-\hat{P})X$, the gradient is
However, there is a more efficient method available based on the reductions approach of agarwal_reductions_2018. In this approach, we use cost minimization loss, which punishes misalignment between $\hat{P}_{i}$ and $Y_{i}$ with a cost and minimizes total costs (zadrozny_cost-sensitive_2003). The cost-sensitive loss function with the standard erroneous classification costs of 1 for all observations, which has the same solution as the standard binary cross entropy loss, can be added to the true positive rate disparity unfairness penalty terms to get the cost minimization penalized loss function:
$L=\sum_{i=1}^{n}\left(\hat{P}_{i}(1-Y_{i})+(1-\hat{P}_{i})Y_{i}\right)+\sum_{g\in G}\lambda_{g}n_{g}\left(\frac{\sum_{k\in r}Y_{k}\hat{P}_{k}}{\sum_{k\in r}Y_{k}}-\frac{\sum_{j\in g}Y_{j}\hat{P}_{j}}{\sum_{j\in g}Y_{j}}\right).$
$L=\sum_{i=1}^{n}\left(\hat{P}_{i}\left(1-Y_{i}+Y_{i}\sum_{g\in G}\lambda_{g}n_{g}\left(\frac{\mathbb{I}_{ri}}{\sum_{k\in r}Y_{k}}-\frac{\mathbb{I}_{gi}}{\sum_{j\in g}Y_{j}}\right)\right)+(1-\hat{P}_{i})Y_{i}\right).$
With this formulation, we extend a theorem from zadrozny_cost-sensitive_2003 to rewrite the cost-sensitive classification problem as a weighted non-cost-sensitive classification problem. In this formulation, the weights are
$W_{i}=|C_{i}^{0}-C_{i}^{1}|=\left|Y_{i}\left(2-\sum_{g\in G}\lambda_{g}n_{g}\left(\frac{\mathbb{I}_{ri}}{\sum_{k\in r}Y_{k}}-\frac{\mathbb{I}_{gi}}{\sum_{j\in g}Y_{j}}\right)\right)-1\right|,$
@@ 3.2 Selecting Penalty Coefficients
Our penalty weight search is a random search, which yields higher maximum scores more efficiently than a grid search (bergstra_random_2012). All the group penalty weight candidates are sampled from identical independent log uniform distributions. We favor this approach over greedy sequential approaches because it allows for optimization with respect to multiple score functions simultaneously. The breadth of the penalty weight candidate search space and number of candidate draws will vary for each fair penalized regression problem, but we include some empirical demonstrations of different choices in our simulation study.
@@ 4 SIMULATION STUDY
Our simulation was designed such that the groups have different data generating processes and unfair true positive rate disparities with respect to the reference group. Let $g_{A}$, $g_{B}$, and $g_{C}$ be the three different groups whose membership was reflected by indicators $\mathbb{I}_{g_{A}}$, $\mathbb{I}_{g_{B}}$, and $\mathbb{I}_{g_{C}}$, respectively. We defined the reference group to be the set of individuals who were not in any of the three groups. The data were generated as follows with $n=1,000,000$:
$X_{1},X_{2}\sim\mathcal{N}(30,15)$
$X_{3},X_{4}\sim Pois(15)$
$X_{5},\dots,X_{9}\sim Bern(p),p=[0.2,0.4,0.6,0.8,0.95]$
$\mathbb{I}_{g_{A}}\sim Bern(\min(\max(c_{0}(X_{9}+0.1)+c_{3}X_{6}+\mathcal{N}(0,0.02),0),1))$
$\mathbb{I}_{g_{B}}\sim Bern(\min(\max(c_{1}(X_{7}+0.05)+c_{4}X_{6}+\mathcal{N}(0,0.02),0),1))$
$\mathbb{I}_{g_{C}}\sim Bern(\min(\max(c_{2}(X_{8}+0.4)+c_{5}X_{6}+\mathcal{N}(0,0.02),0),1))$
$Y$ $\sim Bern\Bigg(\Phi^{-1}\bigg(\frac{X_{1}}{30}-\frac{X_{2}}{15}+\frac{3X_{3}}{50}-\frac{X_{4}}{25}+\frac{1}{100}\Big(2^{X_{5}+X_{6}}+6X_{7}$
[…]

### v2

@@ Abstract
Fair regression methods have the potential to mitigate societal bias concerns in health care, but there has been little work on penalized fair regression when multiple groups experience such bias. We propose a general regression framework that addresses this gap with unfairness penalties for multiple groups. Our approach is demonstrated for binary outcomes with true positive rate disparity penalties and can be efficiently implemented through reduction to a cost-sensitive classification problem. We additionally introduce novel score functions for automatically selecting penalty weights. Our penalized fair regression methods are empirically studied in simulations, where they achieve a fairness-accuracy frontier beyond that of existing comparison methods. Finally, we apply these methods to a national multi-site primary care study of chronic kidney disease to develop a fair classifier for end-stage renal disease. There we find substantial improvements in fairness for multiple race and ethnicity groups who experience societal bias in the health care system without any appreciable loss in overall fit.
Short Running Head: Penalized Fair Regression for Multiple Groups
Keywords: algorithmic bias, chronic kidney disease, fair regression, penalized regression
@@ 1 INTRODUCTION
Fair regression methods modify the loss function to encode fairness measures. These loss functions typically include a single penalty or constraint based on a measure of unfairness for one group (Calders et al., 2013; Berk et al., 2017; Corbett-Davies et al., 2023). Subsequent work has developed additional technical methods for solving single group penalized or constrained fair regression problems (Dwork et al., 2017; Chzhen et al., 2020; Lohaus et al., 2020; Mishler and Kennedy, 2021) and tailored them for applications in cyberbullying (Gencoglu, 2021) and health policy (Zink and Rose, 2020).
@@
Although there is a mathematical equivalence between penalized and constrained methods such that each penalized regression has a corresponding constrained regression with the same solution, most methods formulate their fair regressions as constrained optimization problems. This overlooks that penalty-based approaches have several advantages relative to constraint-based approaches that warrant further development. For one, the approaches are conceptually distinct. A constraint is a user-imposed maximum level of allowable unfairness, while a penalty is a user-imposed tradeoff between accuracy and fairness. These conceptions are appropriate for different applications and their varied social and regulatory contexts. The two approaches may also require divergent computational implementations. Much of the constrained optimization literature involves development of original solution algorithms (Kearns et al., 2018; Komiyama et al., 2018), whereas there are many existing algorithms for minimizing penalized loss functions under different conditions, including gradient descent, Newton’s method, Nelder-Mead, and their variations (Nelder and Mead, 1965; Shanno, 1970; Loshchilov and Hutter, 2019). Outside of the fairness literature, penalized methods are commonly used in statistics applications (Hoerl and Kennard, 1970; Tibshirani, 1996) and fair penalty methods can benefit from advances established in those applications.
@@
Fair regression has the potential to reduce the impact of societal bias in health care. However, only a small number of fair regression works have examined realistic use cases in health care (Zink and Rose, 2020; McGuire et al., 2021; Pfohl et al., 2021, 2022; Wang and Yang, 2025). Health care applications have complex ethical considerations, many of which are unique due to the complexity of medical systems (Chen et al., 2021, 2023; Feng et al., 2025). For example, questions about how to handle certain patient characteristics in clinical algorithms, such as race, are nuanced such that not all fair regression approaches may be appropriate (Cusick et al., 2024; Zink et al., 2024).
@@
A major limitation of most fair regression approaches is that they focus on fairness with respect to a single group when multiple groups typically face unfairness in real world settings (Yang et al., 2020; Chen et al., 2021; Liu et al., 2022; Wang and Yang, 2025). The joint consideration of multiple groups in fair regression is non-trivial, with challenges related to the optimal weighting of the groups as well as the additional computational complexity (Kearns et al., 2018; Narasimhan et al., 2020; Diana et al., 2021). We refer here to multiple groups rather than sensitive attributes to avoid the particular social and sometimes legal connotations of the latter, reflecting the breadth of possible groups for which these methods may be appropriate (Deho et al., 2023).
@@
While limited compared to methods for single groups, there is a small literature on fair constrained regression for multiple groups establishing basic frameworks and efficient solution methods (Kearns et al., 2018; Narasimhan, 2018; Agarwal et al., 2019; Zafar et al., 2019; Narasimhan et al., 2020; Yang et al., 2020; Diana et al., 2021; Hu et al., 2024; Wang and Yang, 2025). Of particular note, Agarwal et al. (2018) proposed a method that reduces fair binary classification to a sequence of cost-sensitive classification problems, enabling the use of efficient off-the-shelf classifiers for fair constrained regression. This approach is flexible across common fairness metrics and was extended in multiple later papers, including to continuous outcomes (Agarwal et al., 2019) and to non-linear but convex constraints, including demographic parity and equalized odds, and convex or fractional-convex losses (Narasimhan, 2018). Much of the constrained optimization literature involves solving a min–max saddle-point game between a predictor and an adversary enforcing fairness constraints, often leading to non-convex objectives, instability, substantial hyperparameter sensitivity, and high computational cost, especially when group sizes are imbalanced.
@@
In contrast, we have identified even fewer papers that implement fair penalized regression with respect to multiple groups, further highlighting the lesser focus on this formulation of the estimation problem. Pérez-Suay et al. (2017) used the Hilbert Schmidt independence criterion for fairness penalties. This method is only compatible with one measure of fairness, demographic parity, which is satisfied when estimated outcomes are independent of groups. A ridge penalty method that builds on Komiyama et al. (2018) to enforce demographic parity was developed in Scutari et al. (2022). Their approach notably includes group identifiers as predictors, which will not be appropriate for some estimation problems. They also proposed an extension to other fairness metrics, including equalized odds. The focus of these first two papers on demographic parity restricts applications in health care, where the most pressing fairness concern may differ from and be incompatible with demographic parity (Seyyed-Kalantari et al., 2021; Chen et al., 2023). Pfohl et al. (2021) and Pfohl et al. (2022) implement fair penalized regression with a single penalty term that measures unfairness with respect to a single categorical group, capturing a subset of multiple group fair regression problems. Lastly, Liu et al. (2022) introduced a fairness measure built on cross-covariance operators on reproducing kernel Hilbert Spaces which were used to construct unfairness penalties. While the approach is flexible with respect to fairness measures, its complexity presents a potential barrier to its implementation and no code is available.
@@
Our application creates a fair penalized regressor classifying whether high-risk individuals with chronic kidney disease (CKD) will develop end stage renal disease (ESRD) in the next year. CKD is prevalent, affecting an estimated 14% of the US adult population (CDC, 2025). In severe cases, CKD progresses to ESRD, requiring either dialysis or kidney transplant, but progress can be delayed through treatment and management. Black, Hispanic, Native American, and Asian patients all have higher incidence and prevalence of ESRD than white patients, and they have worse kidney function, on average, than white patients when initiating treatment for ESRD (Johansen et al., 2025). Identifying patients at risk for ESRD could help these patients engage in more intensive treatment and slow CKD progression, and fair regression is a potential approach to partially address disparities in progression to ESRD. Another key algorithm for defining CKD progression, the estimated glomerular filtration rate (eGFR) equation, was recently revised to remove race adjustment, so CKD may be a context where it is especially appropriate to use a fair regression method that can increase fairness without using race adjustments (Cusick et al., 2024).
@@ 2.1 Fairness Formulation
We use true positive rate equality, based on equal opportunity (Hardt et al., 2016), which holds if the true positive rate of our estimator for $Y$ is the same for members of group $g$ as it is for a reference group $r$. A focus on true positive rate fairness (or, equivalently, false negative rate fairness) is appropriate for the problem of low rates of screening and diagnosis among high-risk patients in certain groups (Seyyed-Kalantari et al., 2021). This is especially true in cases where the benefit of a diagnosis is much greater than the potential harm, as this implies that false negatives are much more costly than false positives and should be the primary fairness concern. We can express different versions of this fairness measure for estimated outcomes and for estimated probabilities. The estimated outcome formulation is more interpretable. However, the estimated probability formulation is differentiable, so it can be easily minimized, and it does not depend on a particular classification threshold.
@@
where $\hat{Y}$ represents predicted outcomes. For estimated probabilities, we substitute $\hat{P}$ for $\hat{Y}$ as has been done in prior literature (Wang and Yang, 2025). This version of the unfairness measure, which we call $U_{EOP}$, is given by:
@@ 2.2 Fairness Measures for Multiple Groups
There is no consensus in the literature on optimal ways to evaluate fairness for multiple groups simultaneously, as much of the past fair regression work characterizes unfairness for each group individually (Agarwal et al., 2018; Komiyama et al., 2018). Given the unfairness metrics in Equation 1 and Equation 2 for a given group and reference, which we will denote $U_{g}$, we study three approaches to aggregating across groups.
@@
    Additionally, we study a maximum across all group-level unfairness measures, based on Ghosh et al. (2022): $U_{Max}=\max_{g\in G}U_{g}$.
@@ 3 METHODS
Our novel penalized fair regression approach generalizes to various fairness metrics, outcome variable types, and estimator architectures. We focus our main demonstration of the method on logistic regression with true positive rate disparity penalties. Web Appendix A contains loss functions for continuous outcomes with the mean residual difference as the unfairness metric.
@@ 3.1 Regression with Fairness Penalties
$L=-\sum_{i=1}^{n}\{Y_{i}\ln(\hat{P}_{i})+(1-Y_{i})\ln(1-\hat{P}_{i})\}+\sum_{g\in G}\lambda_{g}n_{g}\left(\frac{\sum_{k\in r}Y_{k}\hat{P}_{k}}{\sum_{k\in r}Y_{k}}-\frac{\sum_{j\in g}Y_{j}\hat{P}_{j}}{\sum_{j\in g}Y_{j}}\right).$
The above loss function is convex but does not have a closed form minimizing solution. We employ an efficient optimization method based on the reductions approach of Agarwal et al. (2018). In this approach, we use cost minimization loss to reconceptualize the penalized fair regression as a cost-sensitive classification problem, which can be implemented as a weighted non-cost-sensitive classification problem. In this formulation, the weights are
$W_{i}=|C_{i}^{0}-C_{i}^{1}|=\left|Y_{i}\left\{2-\sum_{g\in G}\lambda_{g}n_{g}\left(\frac{\mathbb{I}_{ri}}{\sum_{k\in r}Y_{k}}-\frac{\mathbb{I}_{gi}}{\sum_{j\in g}Y_{j}}\right)\right\}-1\right|,$
and modified outcomes are $Y_{i}^{\prime}=\mathbb{I}(C_{i}^{0}>C_{i}^{1})$. It is then possible to perform the desired fair penalized regression with a wide array of standard classifiers (e.g., logistic regression, support vector machines, decision trees) that are compatible with individual-level sample weighting. More detailed derivations are in Web Appendix B.
@@ 3.2 Selecting Penalty Coefficients
Our penalty weight search is a random search, which yields higher maximum scores more efficiently than a grid search (Bergstra and Bengio, 2012). All the group penalty weight candidates are sampled from identical independent log uniform distributions. We favor this approach over greedy sequential approaches because it allows for optimization with respect to multiple score functions simultaneously. The breadth of the penalty weight candidate search space and number of candidate draws will vary for each fair penalized regression problem, but we include some empirical demonstrations of different choices in our simulation study.
@@ 4 SIMULATION STUDY
Our simulation was designed such that the groups have separate data generating processes and unfair true positive rate disparities with respect to the reference group. We had a population of 1,000,000 per simulation with three groups, with substantial overlap between the groups in all settings. Simulation setting 1 has all moderately sized groups, simulation setting 2 has one small group, and all groups are small in simulation setting 3. The data generating process is available in Web Appendix C.